the perimeter of a rectangle is 36 feet. the length is 2 feet more than three times the width. find the length and width of the rectangle
Answer by Hawksfan(61)

P = 2L + 2W
L = 3W + 2
P = 36
Substitute L into the P equation:
36 = 2(3W + 2) + 2W
36 = 6W +4 +2W
36 = 8W +4
32 = 8W
W = 4
L = 3*4 +2 = 14